OpenStudy (anonymous):

how would i solve this? Given the function g(x) = 4x − 5, find g(−3).

OpenStudy (anonymous):

just replace for x=-3 g(-3)=4*(-3)-5

OpenStudy (anonymous):

Where ever there is an x put -3 in its place. Like so g(x) = 4(-3) - 5 Now solve for g(x) g(x) = -12 - 5 I will let you finish

OpenStudy (anonymous):

-17 but the guy above you subsituted x by g whichyou didnt

OpenStudy (anonymous):

He put -3 in g(x) is that what you are talking about? where he g(-3) ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

That is the correct form. You do not have to do that but that is for showing that x is an input. g(x) is a function and functions take input and you show the input by g(-3), which he is saying, he is inputing -3 for x or x = -3

OpenStudy (anonymous):

thank you! :)

OpenStudy (anonymous):

YW